OpenStudy (anonymous):

\[Determine: \int\limits_{}^{}e^xcosxdx\] Through repeated partial integration.

OpenStudy (experimentx):

use integration by parts, u is cos(x) and v is e^x

sam (.sam.):

\[u = e^x ~~and~~ dv = cosx\] \[du = e^x dx~~ and~~ v = sinx\] \[∫u dv = uv - ∫v du\] \[= e^x sinx - ∫e^x sinx dx\] integration by parts again, \[u = e^x ~~and~~ dv = sin~x\] \[du = e^x dx~~ and ~~v = -cosx\] \[= e^x sinx - [-e^x cosx - ∫-e^x cosx dx]\] \[= e^x sinx - [-e^x cosx + ∫e^x cosx dx]\] \[= e^x sinx + e^x cosx - ∫e^x cosx dx\] if you rewrite it like this, \[∫e^x cosx dx = e^x sinx + e^x cosx - ∫e^x cosx dx\] you can see that that there are\[ ∫e^x cosx dx\] on both side, so, just add \[∫e^x cosx dx\] to both sides -------------------------------------------------- so you got \[2∫e^x cosx = e^x sinx + e^x cosx\] \[∫e^x cosx = \frac{1}{2}(e^x sinx + e^x cosx) +constant\]

OpenStudy (experimentx):

A bit more general approach http://answers.yahoo.com/question/index?qid=20080723161756AA9xWm6

OpenStudy (anonymous):

Thanks! The cycling really circulated my mind, haha :s Glad it had a relatively easy solve. :)